# Incremental SWAP Operator for Entanglement Entropy: Application for Exponential Observables in Quantum Monte Carlo Simulation

### v1

@@ Abstract
We propose a new method to efficiently compute the entanglement entropy (EE) of quantum many-body systems. Our approach, called the incremental SWAP operator method, combines the simplicity of the SWAP operator used in projector quantum Monte Carlo simulations [1] with recent advances in precisely computing exponential observables using incremental algorithms [2, 3]. We apply this technique to obtain accurate EE data at reduced computational cost for 1d and 2d antiferromagnetic Heisenberg models with different bipartition schemes. Using the computed EE data, we extract the area law coefficient, universal logarithmic corrections from Goldstone modes, and the geometric constant, finding quantitative agreement with analytical predictions. Moreover, for the first time in an unbiased numerical simulation of 2d antiferromagnetic Heisenberg model, we successfully obtain reliable universal logarithmic corrections from sharp corners that match expected theoretical values. The consistency between our numerical results and theoretical calculations demonstrates the power of our approach for accessing challenging universal entanglement properties. The extensions of our method to other quantum spin/boson models and the interacting fermion models, are outlined.
@@ I Introduction
where $\rho_{s}$ denotes the spin stiffness at thermodynamic limit, $v$ represents the velocity of Goldstone modes, and $\gamma_{\text{ord}}$ is a universal geometric constant depending on the partition scheme  [^1] Here, we choose to explicitly include $v$ and $\rho_{s}$ because their finite-size corrections are needed in the fitting in Eq. (14).. Especially, the logarithmic correction contains two distinct contributions. $s_{G}=\frac{n_{G}}{2}$ is independent of the shape of $A$ and counts the number of Goldstone modes, it has been numerically computed with successes in Refs. [22, 11, 23, 13]. The other term, $s_{c}$ is the corner correction similar with that in CFT, and it is much smaller in amplitude than $s_{G}$ (with the opposite sign), its accurately extraction is more challenging, with a few less accurate numerical attempts [24, 25, 26, 27].
@@
In fact, early QMC attempts has already been applied to calculate the 2nd Rényi EE in 2d quantum models, mainly by the SWAP operator – an instance of the reduced density matrix for a QMC configuration – based algorithm and its variations [1, 28, 19, 29, 30, 31, 32]. However, directly applying these QMC methods to calculate EE has been shown to have serious issue with convergence [1] (we refer to as the non-convergence problem of SWAP operator hereafter). The underlying cause of this problem has not been clearly recognized or identified. Recently, with the fast developments of the incremental algorithm [33, 34], the non-convergence problem of the EE computation has been greatly reduced (although still not in the SWAP operator calculation), the more reliable data to extract the universal properties of EE from large-scale QMC calculations in 2d bosonic and fermionic system have been obtained [7, 35, 8, 2, 9, 10, 11, 36, 12, 37, 3, 13], and the true reason behind the non-convergence problem begins to be understood.
In this paper, by combining the simplicity of the SWAP operator in the projector quantum Monte Carlo (PQMC) simulation [1, 19] and the recent understanding of computation of the exponential observables in the incremental algorithm [2, 3], we quantitatively demonstrate that the origin of the non-convergence problem of previous EE calculation lies in the fact that EE is an exponential observable, with exponential explosion of coefficient of variation as $L$ increases. We further provide a solution to this problem, dubbed incremental SWAP operator method with improved propagating update scheme for PQMC, to obtain the 2nd Rényi EE with less computational cost and more accurate data. We apply the method upon 1d and 2d antitferromagnetic Heisenberg models and obtain results with better quality than the simple SWAP results, and in the 2d case, we successfully obtained the area law coefficient, the universal logarithmic corrections, both $s_{G}$ and $s_{c}$, and the the geometric constant $\gamma_{\text{ord}}$ in Eq. (2), all of which, are well consistent with the theoretical predictions. We note that previous QMC results have never obtained all the coefficients accurately as we have done here. The extensions of our method to other quantum spin/boson models and the interacting fermion models, are straightforward.
@@ II Non-convergence Problem for EE computation
in terms of the updating weight $W_{C}=w_{l}w_{r}\left\langle V_{l}^{0}\left|\prod\limits_{i=1}^{2m}H(p_{i})\right|V_{r}^{0}\right\rangle$, in other word,
@@
The $\langle\text{SWAP}_{A}\rangle=e^{-S_{A}^{(2)}}$ is an exponential observable [3], and as $L$ increases, $S_{A}^{(2)}\propto L$ according to the area law, causing $e^{-S_{A}^{(2)}}$ to decrease exponentially, which makes estimating $\langle\text{SWAP}_{A}\rangle$ via PQMC extraordinarily challenging for large $L$. And this the reason of the non-convergence problem in the previous SWAP operator calculation of EE. Below, we quantitatively elucidate such problem with numerical data. As depicted in Fig. 2, we plot the QMC sampling distribution and the corresponding histogram of a $10\times 10$ lattice of 2d antiferromagnetic Heisenberg model with “stripe” subsystem for both $\text{SWAP}_{A}$ and $\ln{(\text{SWAP}_{A})}$. Evidently, as shown in Fig. 2 (a) and (b), the vast majority of $\text{SWAP}_{A}$ samples lie extremely close to 0, markedly deviating from a normal distribution, which one can sample well in the QMC simulations. Strikingly, as shown in Fig. 2 (c) and (d), the distribution of $\ln(\text{SWAP}_{A})$ perfectly conforms to normality.
@@
To address such non-convergence problem, we realize $\text{CV}[(\text{SWAP}_{A})^{1/n}]=\sqrt{e^{\sigma^{2}/n^{2}}-1}$, with $n$ to appropriately scale with $L$, can be used to suppress the exponential growth of coefficient of variation. Here, we choose $n=|\langle\ln(\text{SWAP}_{A})\rangle|$, then $\text{CV}[(\text{SWAP}_{A})^{1/n}]$ would decay with $L$, enabling accurate evaluation. As shown in Fig. 3 (c), $\text{CV}[\ln(\text{SWAP}_{A})]$ and $\text{CV}[(\text{SWAP}_{A})^{1/n}]$ indeed exhibit similar decaying behavior as $L$ increases. This proves $(\text{SWAP}_{A})^{1/n}$ can be computed accurately in analogous to $\ln(\text{SWAP}_{A})$, and the sampling of it will render normal distributions.
Figure 3: Coefficient of variation of $\langle\text{SWAP}_{A}\rangle$ and its solution. (a) and (b) show standard deviation $\sigma$ and expectation $\mu$ of $\ln(\text{SWAP}_{A})$ for a bipartition of square lattice Heisenberg model with smooth boundary as a function of system size $L$ in the log-log scale. Both $\sigma$ and $\mu$ of $\ln(\text{SWAP}_{A})$ increase with $L$. The data can be fit with a perfect power law as denoted by the dashed lines. (c) shows $L$ dependence of coefficient of variation, CV, for three observables. CV[$\text{SWAP}_{A}$] increases exponentially while the others $\ln(\text{SWAP}_{A})$ and $(\text{SWAP}_{A})^{1/n}$ stay constant and even decreases with $L$, meaning that they are the ideal QMC observables with normal distribution.
@@ III Incremental SWAP operator Method
with $Z(0)=\sum_{C}W_{C}$ and $Z({n})=\sum_{C}W_{C}\text{SWAP}_{A}$, naturally satisfies the 2nd Rényi EE definition. Subsequently, each increment is evaluated in parallel per PQMC as follows
@@
where $n$ constitutes the ceiling of $|\mu|$. The new sampling observable is $(\text{SWAP}_{A})^{1/n}$ and the new sampling weight is $W_{C}(\text{SWAP}_{A})^{k/n}$. Our incremental SWAP operator method allows us to evaluate the new observables $(\text{SWAP}_{A})^{1/n}$ for each piece of parallel incremental process with controlled statistical errors, as illustrated in Fig. 4(a). As expected, all pieces of the incremental process should generate reliable samples that approximately follow a normal distribution and have a finite variance. To demonstrate this, we have selected two parallel pieces and plotted their sample distributions and histograms in Fig. 4 (b) and (c).
@@
In total, our incremental SWAP operator method with propagating update scheme for evaluating $S^{(2)}_{A}$ has a superior computational complexity of $O(mP)$ with $P=N$ or $1$ for magnetic or paramagnetic states. The full complexity follows as $O(mPL^{\gamma})$ given the requisite $n=|\ln(\text{SWAP}_{A})|\sim L^{\gamma}$ increments. For Heisenberg model, $|\ln(\text{SWAP}_{A})|=0.23L^{1.5}$ in Fig. 3 (b), thus $\gamma=1.5$. Ascertaining $\gamma$ necessitates analysing $\mu(L)$ and $\sigma(L)$ scaling, costing $O(m)$. However, fitting merely requires a few small $L$. Therefore, this negligible overhead contrasts the total cost of evaluating EE.
@@ Acknowledgements.
X.Z. and Y.Q. acknowledge the support from National Key R&D Program of China (Grant No.2022YFA1403400) and from NSFC (Grant No. 12374144). Z.Y.M. thank Menghan Song, Jiarui Zhao, Meng Cheng and Cenke Xu for valuable discussions on related topics, and acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region of China (Project Nos. 17301721, AoE/P701/20, 17309822, C7037-22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper.

### v2

@@ Abstract
We propose a new method to efficiently compute the entanglement entropy (EE) of quantum many-body systems. Our approach, called the incremental SWAP operator method, combines the simplicity of the SWAP operator used in projector quantum Monte Carlo simulations with recent advances in precisely computing exponential observables using incremental algorithms. We apply this technique to obtain accurate EE data at reduced computational cost for 1d and 2d antiferromagnetic Heisenberg models with different bipartition schemes. Using the computed EE data, we extract the area law coefficient, universal logarithmic corrections from Goldstone modes, and the geometric constant, finding quantitative agreement with analytical predictions. Moreover, for the first time in an unbiased numerical simulation of 2d antiferromagnetic Heisenberg model, we successfully obtain reliable universal logarithmic corrections from sharp corners that match expected theoretical values. The consistency between our numerical results and theoretical calculations demonstrates the power of our approach for accessing challenging universal entanglement properties. The extensions of our method to other quantum spin/boson models and the interacting fermion models, are outlined.
@@ I Introduction
where $\rho_{s}$ denotes the spin stiffness at thermodynamic limit, $v$ represents the velocity of Goldstone modes, and $\gamma_{\text{ord}}$ is a universal geometric constant depending on the partition scheme  [^1] Here, we choose to explicitly include $v$ and $\rho_{s}$ because their finite-size corrections are needed in the fitting in Eq. (14).. Especially, the logarithmic correction contains two distinct contributions. $s_{G}=\frac{n_{G}}{2}$ is independent of the shape of $A$ and counts the number of Goldstone modes, which has been numerically computed with success in Refs. [19, 8, 20, 10]. The other term, $s_{c}$ is the corner correction similar with that in CFT, and it is much smaller in amplitude than $s_{G}$ (with the opposite sign), its accurate extraction is more challenging, with a few less accurate numerical attempts [21, 22, 23, 24].
@@
In fact, early QMC attempts have already been applied to calculate the 2nd Rényi EE in 2d quantum models, mainly by the SWAP operator – an instance of the reduced density matrix for a QMC configuration – based algorithm and its variations [27, 28, 16, 29, 30, 31, 32]. However, directly applying these QMC methods to calculate EE has been shown to have serious issue with convergence [27] (we refer to as the non-convergence problem of SWAP operator hereafter). The underlying cause of this problem has not been clearly recognized or identified. Recently, with the fast developments of the incremental algorithm [33, 34], the non-convergence problem of the EE computation has been greatly reduced (although still not in the SWAP operator calculation), the more reliable data to extract the universal properties of EE from large-scale QMC calculations in 2d bosonic and fermionic system have been obtained [4, 35, 5, 25, 6, 7, 8, 36, 9, 37, 26, 10], and the true reason behind the non-convergence problem begins to be understood.
In this paper, by combining the simplicity of the SWAP operator in the projector quantum Monte Carlo (PQMC) simulation [27, 16] and the recent understanding of computation of the exponential observables in the incremental algorithm [25, 26], we quantitatively demonstrate that the origin of the non-convergence problem of previous EE calculation lies in the fact that EE is an exponential observable, with exponential explosion of coefficient of variation as $L$ increases. We further provide a solution to this problem, dubbed incremental SWAP operator method with improved propagating update scheme for PQMC, to obtain the 2nd Rényi EE with less computational cost and more accurate data. We apply the method upon 1d and 2d antiferromagnetic Heisenberg models and obtain results with better quality than the simple SWAP results, and in the 2d case, we successfully obtained the area law coefficient, the universal logarithmic corrections, both $s_{G}$ and $s_{c}$, and the the geometric constant $\gamma_{\text{ord}}$ in Eq. (2), all of which, are well consistent with the theoretical predictions. We note that previous QMC results have never obtained all the coefficients accurately as we have done here. The extensions of our method to other quantum spin/boson models and the interacting fermion models, are straightforward.
@@ II Non-convergence Problem for EE computation
in terms of the updating weight $W_{C}=w_{l}w_{r}\left\langle V_{l}^{0}\left|\prod\limits_{i=1}^{2m}H(p_{i})\right|V_{r}^{0}\right\rangle$; in other words,
@@
The $\langle\text{SWAP}_{A}\rangle=e^{-S_{A}^{(2)}}$ is an exponential observable [26], and as $L$ increases, $S_{A}^{(2)}\propto L$ according to the area law, causing $e^{-S_{A}^{(2)}}$ to decrease exponentially, which makes estimating $\langle\text{SWAP}_{A}\rangle$ via PQMC extraordinarily challenging for large $L$. And this is the reason of the non-convergence problem in the previous SWAP operator calculation of EE. Below, we quantitatively elucidate such problem with numerical data. As depicted in Fig. 2, we plot the QMC sampling distribution and the corresponding histogram of a $10\times 10$ lattice of 2d antiferromagnetic Heisenberg model with “stripe” subsystem for both $\text{SWAP}_{A}$ and $\ln{(\text{SWAP}_{A})}$. Evidently, as shown in Fig. 2 (a) and (b), the vast majority of $\text{SWAP}_{A}$ samples lie extremely close to 0, markedly deviating from a normal distribution, which one can sample well in the QMC simulations. Strikingly, as shown in Fig. 2 (c) and (d), the distribution of $\ln(\text{SWAP}_{A})$ perfectly conforms to normality.
@@
To address such non-convergence problem, we realize $\text{CV}[(\text{SWAP}_{A})^{1/n}]=\sqrt{e^{\sigma^{2}/n^{2}}-1}$, with $n$ to appropriately scale with $L$, which can be used to suppress the exponential growth of coefficient of variation. Here, we choose $n$ as the ceiling integer of $|\langle\ln(\text{SWAP}_{A})\rangle|$, then $\text{CV}[(\text{SWAP}_{A})^{1/n}]$ would decay with $L$, enabling accurate evaluation. In practice, choosing an integer $n$ such that $\text{CV}[(\text{SWAP}_{A})^{1/n}]$decreases or approaches a constant with $L$ is sufficient to guarantee the convergence. A smaller CV means that we just need fewer samples to calculate entanglement entropy accurately. As shown in Fig. 3 (c), $\text{CV}[\ln(\text{SWAP}_{A})]$ and $\text{CV}[(\text{SWAP}_{A})^{1/n}]$ indeed exhibit similar decaying behavior as $L$ increases. This proves $(\text{SWAP}_{A})^{1/n}$ can be computed accurately in analogous to $\ln(\text{SWAP}_{A})$, and the sampling of it will render normal distributions.
Figure 3: Coefficient of variation of $\langle\text{SWAP}_{A}\rangle$ and its solution. (a) and (b) show standard deviation $\sigma$ and expectation $\mu$ of $\ln(\text{SWAP}_{A})$ for a bipartition of square lattice Heisenberg model with smooth boundary as a function of system size $L$ in the log-log scale. Both $\sigma$ and $\mu$ of $\ln(\text{SWAP}_{A})$ increase with $L$. The data can be fit with a perfect power law as denoted by the dashed lines. (c) shows $L$ dependence of coefficient of variation, CV, for three observables. Here, we choose $n$ as the ceiling integer of $|\langle\ln(\text{SWAP}_{A})\rangle|$. CV[$\text{SWAP}_{A}$] increases exponentially while the others $\ln(\text{SWAP}_{A})$ and $(\text{SWAP}_{A})^{1/n}$ stay constant and even decreases with $L$, meaning that they are the ideal QMC observables with normal distribution.
@@ III Incremental SWAP operator Method
with $Z(0)=\sum_{C}W_{C}$ and $Z({n})=\sum_{C}W_{C}\text{SWAP}_{A}$, naturally satisfying the 2nd Rényi EE definition. Subsequently, each increment is evaluated in parallel per PQMC as follows
@@
where $n$ constitutes the ceiling integer of $|\mu|$. The new sampling observable is $(\text{SWAP}_{A})^{1/n}$ and the new sampling weight is $W_{C}(\text{SWAP}_{A})^{k/n}$. Our incremental SWAP operator method allows us to evaluate the new observables $(\text{SWAP}_{A})^{1/n}$ for each piece of parallel incremental process with controlled statistical errors, as illustrated in Fig. 4(a). As expected, all pieces of the incremental process should generate reliable samples that approximately follow a normal distribution and have a finite variance. To demonstrate this, we have selected two parallel pieces and plotted their sample distributions and histograms in Fig. 4 (b) and (c).
It is necessary to discuss the effects of the accumulated error here. we simulated each factor in Eq.  (10) independently, which will give rise to a standard deviation factor $\sigma_{i}=e^{\mu/n^{2}+\sigma^{2}/2n^{2}}\sqrt{e^{\sigma^{2}/n^{2}}-1}$, and $\mu_{i}=e^{\mu/n^{2}+\sigma^{2}/2n^{2}}$. According to the error propagation equation, the accumulated error for the product of $n$ factors in Eq.  (10) could be demonstrated with coefficient of variation $\text{CV}_{total}=\sqrt{\sum_{i}^{n}\sigma_{i}^{2}/{\mu_{i}^{2}}}\sim\sigma/\sqrt{|\mu|}$. It is crucial to realize that our incremental SWAP operator method successfully reduces the total coefficient of variation from exponential growth to polynomial growth as $L$ increases.
@@
In total, our incremental SWAP operator method with propagating update scheme for evaluating $S^{(2)}_{A}$ has a superior computational complexity of $O(mP)$ with $P=N$ or $1$ for magnetic or paramagnetic states. The full complexity follows as $O(mPL^{\gamma})$ given the requisite $n=|\ln(\text{SWAP}_{A})|\sim L^{\gamma}$ increments. For Heisenberg model, $|\ln(\text{SWAP}_{A})|=0.23L^{1.5}$ in Fig. 3 (b), thus $\gamma=1.5$. Ascertaining $\gamma$ necessitates analysing $\mu(L)$ and $\sigma(L)$ scaling, costing $O(m)$. However, fitting merely requires a few small $L$. Therefore, this negligible overhead contrasts the total cost of evaluating EE. Up to now, the popular method to measure entanglement entropy with high efficiency and precision is non-equilibrium incremental algorithm [34, 8]. The computational complexity of this method in PQMC is $m\times n_{qt}$, where $m$ is the projector length and $n_{qt}$ is the quenching time. And such incremental algorithm can overcome the convergence issue of entanglement entropy computation when the quenching time $n_{qt}$ is sufficiently long. However, a quantitative value for $n_{qt}$ has not been determined yet, while the advantage of our method is that the convergence of the calculation can be guaranteed by determining the total increments $n$ precisely.
@@ Acknowledgements.
X.Z. and Y.Q. acknowledge the support from National Key R&D Program of China (Grant No.2022YFA1403400) and from NSFC (Grant No. 12374144). Z.Y.M. and Y.D.L. acknowledge the support from the Research Grants Council (RGC) of Hong Kong Special Administrative Region of China (Project Nos. 17301721, AoE/P701/20, 17309822, C7037-22GF, 17302223), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22), the GD-NSF (no.2022A1515011007) and the HKU Seed Funding for Strategic Interdisciplinary Research. We thank the Beijng PARATERA Tech CO.,Ltd. (URL: https://cloud.paratera.com) for providing HPC resources that have contributed to the research results reported within this paper.